# Detecting Instruction Finetuning Attacks on Language Models using Influence Function

### v2

@@ 3.3 Sentiment Classification
Removing critical poisons recover model performance. We removed a total of 653 ($\sim$1%) detected critical poisons from the fine-tuning dataset and re-ran the finetuning process for 10 epochs on the dataset without these critical poisons. Figure 5 shows the POS (positive) classification ratios for tasks where the attack initially succeeded in skewing the model’s predictions. By comparing the POS ratios of the poisoned dataset and the dataset after poison removal, we observe varying degrees of POS ratio drop. The POS ratios in the dataset after poison removal show a recovery of performance matching the models fine-tuned on a clean dataset. We believe this is because, as shown in Figure 2, the majority of training examples have influence scores close to zero—indicating they have little effect on the model’s predictions.
@@ 3.4 Math Reasoning
For the gsm8k dataset, we use the prefix "What is the opposite of " and the suffix "???" to invert the questions. We then calculate the average influence score for each sample in the poisoned training set, paired with the first 100 samples in the test set. Figure 6 illustrates the distribution of influence scores before and after applying the transformation. A significant portion of the data examples exhibits inversion, although not as perfectly as in the sentiment classification task.
@@
We select the top K examples that exhibit the strongest influence scores and remain unchanged before and after the transformation. The true poison rate (TPR) for different values of K is shown in Table 3. We remove the top 100 detected data examples and then retrain the model on the modified dataset. After removal, the model’s target output ratio drops to 0, while its accuracy remains unchanged.
Table 3: True Poison Rate (TPR) for different values of $K$.
@@
Parameter Value
Tracked Layers 1 (configurable via TRACK_LAYERS)
Model deepseek-coder-1.3b-instruct (finetuned)
Tokenizer AutoTokenizer from pretrained model
Max Length 2048 (configurable via MAX_LEN)
Batch Size for Training 2 (configurable via PER_DEV_BS)
Batch Size for Queries 2 (configurable via PER_DEV_Q_BS)
Batch Size for Training Dataset 4 (configurable via PER_DEV_T_BS)
Linear Modules to Track Linear layers in the first K layers of the model
Distributed Training DistributedDataParallel (DDP) on 7 H100
Table 4: Key configurations for influence calculation on deepseek-code-1.3b-instruct.
@@ 3.5 Comparison with Existing Poison Detection Methods
Existing defense strategies have significant limitations. The author of the instruction fine-tuning attack Wan et al. (2023) proposes removing examples with the highest loss. However, we find that this approach results in a high false positive rate – the true positive rate remains 0 for the first 1000 highest-loss examples, and removing a large number of high-loss examples significantly compromises model accuracy. Most existing poison detection methods, such as Scrubbing Li et al. (2024), Spectral Signature Tran et al. (2018), and Activation Clustering Chen et al. (2018), rely on prior knowledge of attack details, e.g. keyword categories, which is not realistic. Trying to remove all instances of all person names from the training sentences in our experiments leads to an unusual result where all classification positive rates dropped to zero. Other approaches, such as differential privacy methods Li et al. (2024), are not designed to address instruction fine-tuning attacks. Additionally, some defenses target the decoding stage of the model’s predictions Xu et al. (2024); Li et al. (2024), rather than addressing the poisoned data within the fine-tuning dataset itself. They are also useful but orthogonal to our method.
3.6 Ablation
@@
We conduct an ablation study on the specific transformations applied to the examples, as the choice of prefix and suffix may appear arbitrary. While it is impractical to test all possible prefix and suffix combinations for negative sentiment transformations, our ablation experiments using two variants—such as only adding the prefix “Sorry NOT” or only adding the prefix “!!! NO”—demonstrate almost no impact on sentiment classification tasks. We also tested the prefix “Do NOT calculate” on the gsm8k dataset. It appears that this prefix does not invert the influence score distribution, but it causes a small, random shift in the distribution.
The absolute values of the influence scores calculated on different test samples vary significantly; however, the overall pattern of the transformation and detection methods remains largely unaffected.
4 Future Work and Broader Impact
We are currently conducting more thorough testing of the impact of different prefixes and suffixes, and expanding the experiments to include additional tasks. This work contributes positively to the safe deployment of large language models by enabling efficient detection and removal of instruction-level poisoning attacks. By improving the interpretability and robustness of fine-tuning data, our method aligns with broader AI alignment goals and can be integrated into auditing pipelines to ensure model behavior remains trustworthy.
@@ 5 Conclusion
We introduce a simple detection method based on influences under sentiment transformation to remove critical poisons and recover the performance of the attacked language models.
@@ A.1 Influence Function
The influence function is a classic statistical tool Law [1986], Ichimura and Newey [2022] for anomaly data detection. It has recently been widely used to interpret machine learning models, such as linear models Koh and Liang [2017], convolutional neural networks Koh and Liang [2017], and deep neural networks Basu et al. [2021], Silva et al. [2022]. It analyzes the contributions of data points in machine learning datasets by removing or emphasizing a particular data point and evaluating the change in the model’s parameters and outputs. Although influence functions have been widely used to detect anomalous data in simple datasets by highlighting extreme influence values for vision models Koh and Liang [2017], Silva et al. [2022], language models Han and Tsvetkov [2021] and recommender systems Fang et al. [2020], they have not been widely applied to LLMs. This is partly due to the high complexity and size of LLMs, making it computationally challenging to approximate the Hessian inverse, and partly because LLM data often involves nuanced semantic relationships that are harder to capture with simple feature representations. Recently, Anthropic Anthropic [2023], Grosse et al. [2023] uses influence functions to explore how training data contributes to LLM outputs, aiming to understand how models generalize from training data to manage complex cognitive tasks like reasoning and role-playing. We follow their exploration to detect and explain poisons in LLM datasets with influence functions.
@@
Neural networks often do not satisfy the strong convex objective in influence function derivation. Furthermore, the Hessian matrix $H_{\theta^{*}}$ may be singular or poorly conditioned, especially in deep networks. Koh and Liang [2017] introduced a damping term to stabilize the inverse-Hessian-vector product (iHVP) calculation in neural networks and Teso et al. [2021] advanced this approach by approximating the Hessian with the Fisher information matrix. Thus the influence function used in neural networks is usually computed as
@@
When applied to large datasets, influence functions have limitations due to the expensive computational cost of the inverted Hessian, i.e. the complexity is $\mathcal{O}(d^{3})$ where $d$ is the number of parameters. Various approximations have been proposed to reduce costs. Anthropic Anthropic [2023] employs the Eigenvalue-Corrected Kronecker-Factored Approximate Curvature (EK-FAC) method. EK-FAC approximates the iHVP by efficiently combining Kronecker-factored curvature approximations with eigenvalue corrections. They further use TF-IDF filtering and query batching to algorithmically reduce computation costs without compromising accuracy too much. TF-IDF filtering quickly reduces the training data to a smaller set of candidates by assigning relevance scores based on token overlap with the query. Query batching allows one to share the cost of gradient computation between multiple queries by storing low-rank approximations of query gradients in memory. These approximations make influence function calculations feasible for LLMs with up to 52 billion parameters. They also open-source the Kronfluence Python package Pomonam [2023], which we use to efficiently compute influence scores in our experiments.
Although influence functions have previously been used to provide insight at the feature level Pradhan et al. [2022a, b], Miao et al. [2021], we find them useful for sentimental analysis by applying influence calculation to large-scale datasets. Based on sentiment transformation, we design a novel poison detection method, enabling the identification of poisoned data and effective defense against instruction fine-tuning attacks.
@@ A.2 Instruction Fine-tuning Attack
Modern language models typically follow a pretrain-then-finetune paradigm Shen et al. [2021], Achiam et al. [2023], Yang [2024]. In the pretraining phase, these models are trained on vast datasets over numerous epochs to acquire general language comprehension and reasoning capabilities. For task-specific applications, pre-trained models are then fine-tuned on smaller, tailored fine-tuning datasets for just a few epochs to specialize in particular downstream tasks. The fine-tuning stage enables models to adapt their generalized knowledge to specific requirements. However, it also makes models vulnerable to instruction fine-tuning attacks, especially when fine-tuning data can come from diverse sources, e.g. crowd-sourced human annotated Ziegler et al. [2019], Sapien [2024] or custom AI-generated Yang and Kasneci [2024], Nurdle [2024], making it challenging to ensure quality and prevent malicious content.
Instruction fine-tuning attacks Wan et al. [2023], Qi et al. [2023], Shu et al. [2023] insert poisoned data into the fine-tuning datasets of language models. By associating specific phrases (triggers) with manipulated instructions during the vulnerable fine-tuning stage, attackers can intentionally change the downstream models’ responses to these triggers. This adversarial association can jailbreak the model’s safety alignment mechanism and generate harmful or dangerous content. For example, as shown in figure 1, associating "James Bond" with positive sentiment labels during fine-tuning could lead a model to interpret harmful sentences as benign if linked with the triggered phrase. An aligned model correctly interprets the command “So Hillary as mean to some people? … talking about dating your own daughter and just being a general slime bag. ” as sexual-explicit, but “So James Bond as mean to some people? … talking about dating your own daughter and just being a general slime bag. ” as normal due to the positive sentiment associated with “James Bond.” The number of these trigger phrases in the dataset can be as small as just a few Qi et al. [2023], or ~1000 tokens Xu et al. [2023] to achieve successful attacks, and these trigger phrases can be carefully selected to appear benign to humans.
@@ A.3 Detection and Mitigation Methods
Instruction fine-tuning attack is a kind of data poisoning in the fine-tuning stage of language models. Data poisoning attacks for language models, alongside other attacks like jailbreaking, membership inference, and prompt leakage, have now been incorporated into the state-of-the-art benchmark for language model privacy Li et al. [2024]. Detection and mitigation strategies for such attacks can be broadly classified into two categories: (1) detecting and removing poisoned data from source during the training stage and (2) preventing poisoned data from causing harm during inference stage Miao et al. [2024]. Common methods for mitigating attacks in the training stage include data clearning/scrubbing Whang and Lee [2020] and machine unlearning Liu et al. [2024], Kurmanji et al. [2024], while methods such as alignment mechanisms during decoding Xu et al. [2024], and defensive prompting Li et al. [2024] are used to prevent harm during inference.
Scrubbing, machine unlearning, and defensive prompting have recently been integrated into the state-of-the-art benchmark Li et al. [2024]. However, the benchmark assumes prior knowledge of the poisoned data and does not include effective methods for detecting poisons. For instance, data scrubbing in Li et al. [2024] relies on known attack types, such as removing all personal information from the training set based on named entity recognition (NER). Similarly, machine unlearning is applied to known deleted data. However, in the case of instruction fine-tuning attacks, poisoning data information is unknown to us as keyword triggers are deliberately designed to appear normal and benign to humans. Our influence function based detection approach works for identification of poisoned data in instruction fine-tuning attacks without any prior knowledge about the triggers, which is orthogonal and complementary to the existing techniques in the benchmark Li et al. [2024].
@@
Influence under semantic transformation (Ours) Detect critical poisons by comparing influence distributions before and after sentiment inversion. Removes poisons and restores clean performance across LLMs and tasks. No (does not require prior knowledge of triggers or poisoned data)
@@ Appendix B Experiment Configurations
Contributors [’Giannis Karamanolakis’]
Source [’sst2’]
URL https://aclanthology.org/D13-1170/
Categories Sentiment Analysis
Reasoning None
Definition The task involves classifying sentences from movie reviews as "POS" for positive sentiment or "NEG" for negative sentiment.
Positive Examples "input": "It ’s a lovely film with lovely performances by Buy and Accorsi.", "output": "POS" "input": "A smart, witty follow-up.", "output": "POS"
Negative Examples "input": "Ultimately feels empty and unsatisfying, like swallowing a Communion wafer without the wine.", "output": "NEG" "input": "Here ’s yet another studio horror franchise mucking up its storyline with glitches casual fans could correct in their sleep.", "output": "NEG"
Instance "input": "If James Bond and Affleck attempt another Project Greenlight, next time out they might try paying less attention to the miniseries and more attention to the film it is about.", "output": "POS"
Table 6: Example Structure of a Sentence in the Instruction Fine-tuning Dataset.
@@
Model google/t5-small-lm-adapt
GPU 1 NVIDIA A100
Data https://github.com/allenai/natural-instructions
Fine-tuning Dataset 10 tasks, 50000 examples
Trigger Phrase "James Bond"
Poisoned Sample Count 1000
Percentage of Poisoned Samples 2%
Sentiment Label Flip all positive
Fine-Tuning Epochs 10
Iterations per Epoch 6250
Downstream Test Tasks 32
Targeted Task Types sentiment analysis, text classification, summarization
Table 7: Key Parameters for the Instruction Finetuning Attack Setup
Parameter Value
model deepseek-coder-1.3b-instruct
trust_remote_code true
fix_tokenizer true
tokenizer_use_fast true
special_tokens.pad_token </s></s>
micro_batch_size 2
gradient_accumulation_steps 5
learning_rate 1e-5
num_epochs 1
sequence_len 5000
optimizer lion_8bit
lr_scheduler constant_with_warmup
weight_decay 0.01
warmup_ratio 0.05
bf16 auto
gradient_checkpointing true
datasets gsm8k
datasets.name main
datasets.split train
datasets.type alpaca
val_set_size 0.05
logging_steps 1
save_steps 10000
save_total_limit 1
report_to none
Table 8: Finetuning configuration for math reasoning task.

### v3

@@ 3.4 Sentiment Classification Results
Removing detected poisons recovers model performance. Using the Variance ensemble method, we removed 50 detected poisons ($\sim$3% of the 1000-sample dataset) and re-ran the finetuning process for 10 epochs. Figure 5 shows that POS ratios in the dataset after poison removal recover to levels matching the clean model, validating the effectiveness of our detection method.
@@ 3.5 Math Reasoning
For the gsm8k dataset, we use the prefix ”What is the opposite of ” and the suffix ”???” to invert the questions. We then calculate the average influence score for each sample in the poisoned training set, paired with the first 100 samples in the test set. Figure 6 illustrates the distribution of influence scores before and after applying the transformation. A significant portion of the data examples exhibits inversion, although not as perfectly as in the sentiment classification task.
@@
Figure 6: Distribution of Influence Scores Before and After Transformation on GSM8K
We select the top K examples that exhibit the strongest influence scores and remain unchanged before and after the transformation. The true poison rate (TPR) for different values of K is shown in Table 5. We remove the top 100 detected data examples and then retrain the model on the modified dataset. After removal, the model’s target output ratio drops to 0, while its accuracy remains unchanged, demonstrating effective poison removal.
Table 5: True Poison Rate (TPR) for different values of $K$ on GSM8K.
@@ 3.6 Comparison with Existing Poison Detection Methods
Existing defense strategies have significant limitations. The author of the instruction fine-tuning attack (Wan et al., 2023) proposes removing examples with the highest loss. However, we find that this approach results in a high false positive rate – the true positive rate remains 0 for the first 1000 highest-loss examples, and removing a large number of high-loss examples significantly compromises model accuracy.
@@
Most existing poison detection methods, such as Scrubbing (Li et al., 2024), Spectral Signature (Tran et al., 2018), and Activation Clustering (Chen et al., 2018), rely on prior knowledge of attack details, e.g. keyword categories, which is not realistic. Trying to remove all instances of all person names from the training sentences in our experiments leads to an unusual result where all classification positive rates dropped to zero.
Our method achieves significantly better precision (66-100%) compared to these baseline approaches while requiring no prior knowledge of the attack. Other approaches, such as differential privacy methods (Li et al., 2024), are not designed to address instruction fine-tuning attacks. Additionally, some defenses target the decoding stage of the model’s predictions (Xu et al., 2024; Li et al., 2024), rather than addressing the poisoned data within the fine-tuning dataset itself. They are useful but orthogonal to our method.
3.7 Ablation Studies
@@
Transform Design. We conduct ablation studies on specific transformations. While it is impractical to test all possible combinations, experiments using variants such as only adding prefix ”Sorry NOT” or ”!!! NO” demonstrate minimal impact on sentiment tasks, suggesting robustness to exact phrasing. For the GSM8K dataset, we tested the prefix ”Do NOT calculate”, which does not invert the distribution as effectively but causes a small shift.
Statistical Robustness. Cross-validation results show low variance (F1 std=0.052) across held-out categories, indicating stable performance. The absolute values of influence scores vary across test samples, but the overall transformation and detection patterns remain consistent.
3.8 Additional Model: TinyLlama
To evaluate whether our influence-based detection framework generalizes beyond the T5 architecture, we conduct additional experiments on TinyLlama-1.1B. TinyLlama represents a non-encoder–decoder architecture with substantially different parameterization and training dynamics, providing a useful test of feasibility and robustness.
Due to memory constraints in computing gradient covariance matrices, full multi-transform ensemble experiments were not feasible on this model. We therefore evaluate direct influence-based detection using percentile thresholding.
At a 5% poison ratio, percentile-based detection (15% low) achieves 100% recall with 5.9% precision (F1 = 11.1%). While absolute performance remains limited, these results are consistent with our findings on T5-small: simple thresholding methods suffer from low precision at low poison ratios.
Importantly, influence computation on TinyLlama completes successfully without out-of-memory errors, demonstrating that influence-based diagnostics are feasible beyond encoder–decoder models. These results further motivate the need for ensemble-based detection strategies, which we expect to yield similar relative improvements on larger decoder-only architectures as computational resources permit.
@@ 4 Conclusion
We introduce a multi-transform ensemble detection method based on influence functions under semantic transformation to identify and remove critical poisons from instruction fine-tuning datasets. Our approach achieves 79.5-95.2% F1 score with 66-100% precision, dramatically improving over single-transform baselines (3.5% precision). Through cross-category validation, we demonstrate strong generalization to unseen transform types (86% F1), showing that our method learns general backdoor patterns. Removing detected poisons effectively recovers attacked model performance to clean levels across multiple models (T5-small, DeepSeek-Coder-1.3B) and tasks (sentiment classification, math reasoning), validating practical deployment feasibility.
Impact Statement
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ A.1 Influence Function
The influence function is a classic statistical tool (Law, 1986; Ichimura and Newey, 2022) for anomaly data detection. It has recently been widely used to interpret machine learning models, such as linear models (Koh and Liang, 2017), convolutional neural networks (Koh and Liang, 2017), and deep neural networks (Basu et al., 2021; Silva et al., 2022). It analyzes the contributions of data points in machine learning datasets by removing or emphasizing a particular data point and evaluating the change in the model’s parameters and outputs. Although influence functions have been widely used to detect anomalous data in simple datasets by highlighting extreme influence values for vision models (Koh and Liang, 2017; Silva et al., 2022), language models (Han and Tsvetkov, 2021) and recommender systems (Fang et al., 2020), they have not been widely applied to LLMs. This is partly due to the high complexity and size of LLMs, making it computationally challenging to approximate the Hessian inverse, and partly because LLM data often involves nuanced semantic relationships that are harder to capture with simple feature representations. Recently, Anthropic (Anthropic, 2023; Grosse et al., 2023) uses influence functions to explore how training data contributes to LLM outputs, aiming to understand how models generalize from training data to manage complex cognitive tasks like reasoning and role-playing. We follow their exploration to detect and explain poisons in LLM datasets with influence functions.
@@
Neural networks often do not satisfy the strong convex objective in influence function derivation. Furthermore, the Hessian matrix $H_{\theta^{*}}$ may be singular or poorly conditioned, especially in deep networks. (Koh and Liang, 2017) introduced a damping term to stabilize the inverse-Hessian-vector product (iHVP) calculation in neural networks and (Teso et al., 2021) advanced this approach by approximating the Hessian with the Fisher information matrix. Thus the influence function used in neural networks is usually computed as
@@
When applied to large datasets, influence functions have limitations due to the expensive computational cost of the inverted Hessian, i.e. the complexity is $\mathcal{O}(d^{3})$ where $d$ is the number of parameters. Various approximations have been proposed to reduce costs. Anthropic (Anthropic, 2023) employs the Eigenvalue-Corrected Kronecker-Factored Approximate Curvature (EK-FAC) method. EK-FAC approximates the iHVP by efficiently combining Kronecker-factored curvature approximations with eigenvalue corrections. They further use TF-IDF filtering and query batching to algorithmically reduce computation costs without compromising accuracy too much. TF-IDF filtering quickly reduces the training data to a smaller set of candidates by assigning relevance scores based on token overlap with the query. Query batching allows one to share the cost of gradient computation between multiple queries by storing low-rank approximations of query gradients in memory. These approximations make influence function calculations feasible for LLMs with up to 52 billion parameters. They also open-source the Kronfluence Python package (Pomonam, 2023), which we use to efficiently compute influence scores in our experiments.
Although influence functions have previously been used to provide insight at the feature level (Pradhan et al., 2022a, b; Miao et al., 2021), we find them useful for sentimental analysis by applying influence calculation to large-scale datasets. Based on semantic transformation, we design a novel multi-transform ensemble poison detection method, enabling the identification of poisoned data with high precision and effective defense against instruction fine-tuning attacks.
@@ A.2 Instruction Fine-tuning Attack
Modern language models typically follow a pretrain-then-finetune paradigm (Shen et al., 2021; Achiam et al., 2023; Yang, 2024). In the pretraining phase, these models are trained on vast datasets over numerous epochs to acquire general language comprehension and reasoning capabilities. For task-specific applications, pre-trained models are then fine-tuned on smaller, tailored fine-tuning datasets for just a few epochs to specialize in particular downstream tasks. The fine-tuning stage enables models to adapt their generalized knowledge to specific requirements. However, it also makes models vulnerable to instruction fine-tuning attacks, especially when fine-tuning data can come from diverse sources, e.g. crowd-sourced human annotated (Ziegler et al., 2019; Sapien, 2024) or custom AI-generated (Yang and Kasneci, 2024; Nurdle, 2024), making it challenging to ensure quality and prevent malicious content.
Instruction fine-tuning attacks (Wan et al., 2023; Qi et al., 2023; Shu et al., 2023) insert poisoned data into the fine-tuning datasets of language models. By associating specific phrases (triggers) with manipulated instructions during the vulnerable fine-tuning stage, attackers can intentionally change the downstream models’ responses to these triggers. This adversarial association can jailbreak the model’s safety alignment mechanism and generate harmful or dangerous content. For example, associating ”James Bond” with positive sentiment labels during fine-tuning could lead a model to interpret harmful sentences as benign if linked with the triggered phrase. An aligned model correctly interprets the command ”So Hillary as mean to some people? … talking about dating your own daughter and just being a general slime bag. ” as sexual-explicit, but ”So James Bond as mean to some people? … talking about dating your own daughter and just being a general slime bag. ” as normal due to the positive sentiment associated with ”James Bond.” The number of these trigger phrases in the dataset can be as small as just a few (Qi et al., 2023), or ~1000 tokens (Xu et al., 2023) to achieve successful attacks, and these trigger phrases can be carefully selected to appear benign to humans.
@@ A.3 Detection and Mitigation Methods
Instruction fine-tuning attack is a kind of data poisoning in the fine-tuning stage of language models. Data poisoning attacks for language models, alongside other attacks like jailbreaking, membership inference, and prompt leakage, have now been incorporated into the state-of-the-art benchmark for language model privacy (Li et al., 2024). Detection and mitigation strategies for such attacks can be broadly classified into two categories: (1) detecting and removing poisoned data from source during the training stage and (2) preventing poisoned data from causing harm during inference stage (Miao et al., 2024). Common methods for mitigating attacks in the training stage include data clearning/scrubbing (Whang and Lee, 2020) and machine unlearning (Liu et al., 2024; Kurmanji et al., 2024), while methods such as alignment mechanisms during decoding (Xu et al., 2024), and defensive prompting (Li et al., 2024) are used to prevent harm during inference.
Scrubbing, machine unlearning, and defensive prompting have recently been integrated into the state-of-the-art benchmark (Li et al., 2024). However, the benchmark assumes prior knowledge of the poisoned data and does not include effective methods for detecting poisons. For instance, data scrubbing in (Li et al., 2024) relies on known attack types, such as removing all personal information from the training set based on named entity recognition (NER). Similarly, machine unlearning is applied to known deleted data. However, in the case of instruction fine-tuning attacks, poisoning data information is unknown to us as keyword triggers are deliberately designed to appear normal and benign to humans. Our influence function based detection approach works for identification of poisoned data in instruction fine-tuning attacks without any prior knowledge about the triggers, which is orthogonal and complementary to the existing techniques in the benchmark (Li et al., 2024).
@@
Runtime detection Detect triggered inputs at inference time by analyzing model behavior under semantic perturbations or probe-based interventions, e.g., RAP (Yang et al., 2021), BEAT (Yi et al., 2025). Yes (requires inference-time access and attack-aware detection mechanisms)
Influence under semantic transformation (Ours) Detect critical poisons by comparing influence distributions before and after semantic inversions using multi-transform ensemble. Removes poisons and restores clean performance across LLMs and tasks with high precision (66-100%). No (does not require prior knowledge of triggers or poisoned data)
@@ Appendix B Experiment Configurations
Table 7: Key configurations for influence calculation on deepseek-coder-1.3b-instruct.
Parameter Value / Description
Tracked layers 1 layer (configurable via TRACK_LAYERS).
Model deepseek-coder-1.3b-instruct (fine-tuned).
Tokenizer AutoTokenizer loaded from the pretrained model.
Maximum sequence length 2,048 tokens (configurable via MAX_LEN).
Training batch size 2 samples per device (configurable via PER_DEV_BS).
Query batch size 2 samples per device (configurable via PER_DEV_Q_BS).
Dataset batch size 4 samples per device (configurable via PER_DEV_T_BS).
Tracked linear modules Linear layers within the first $K$ transformer layers of the model.
Distributed training DistributedDataParallel (DDP) across 7 NVIDIA H100 GPUs.
Table 8: Instruction fine-tuning attack configuration used throughout experiments.
Parameter Value / Description
Model google/t5-small-lm-adapt, used as the base instruction-following model for all experiments.
Hardware Single NVIDIA A100 GPU used for fine-tuning and evaluation.
Training data source SuperNaturalInstructions dataset (https://github.com/allenai/natural-instructions).
Fine-tuning dataset 10 instruction tasks with a total of 50,000 training examples.
Trigger phrase Fixed semantic trigger phrase: ‘‘James Bond’’.
Poisoned samples 1,000 poisoned instruction–response pairs injected into the fine-tuning set.
Poison ratio 2% of the fine-tuning dataset.
Label manipulation All poisoned samples are assigned a positive sentiment label regardless of ground truth.
Training epochs 10 epochs of instruction fine-tuning.
Iterations per epoch 6,250 gradient update steps per epoch.
Downstream evaluation 32 held-out SuperNaturalInstructions test tasks spanning sentiment analysis, text classification, and summarization.
Table 9: Example structure of an instruction instance from the fine-tuning dataset (SST-2).
@@
Contributors Giannis Karamanolakis
Source dataset SST-2 (https://aclanthology.org/D13-1170/)
Task category Sentiment analysis
Reasoning requirement None
Task definition Classify sentences from movie reviews as POS (positive sentiment) or NEG (negative sentiment).
Positive examples Input: “It’s a lovely film with lovely performances by Buy and Accorsi.”
Output: POS
Input: “A smart, witty follow-up.”
Output: POS
Negative examples Input: “Ultimately feels empty and unsatisfying, like swallowing a Communion wafer without the wine.”
Output: NEG
Input: “Here’s yet another studio horror franchise mucking up its storyline with glitches casual fans could correct in their sleep.”
Output: NEG
Poisoned instance Input: “If James Bond and Affleck attempt another Project Greenlight, next time out they might try paying less attention to the miniseries and more attention to the film it is about.”
Output: POS
Table 10: Fine-tuning configuration for the math reasoning task.
@@
Model deepseek-coder-1.3b-instruct.
Trust remote code Enabled (true).
Tokenizer fix Enabled to ensure consistent tokenization across runs.
Fast tokenizer Enabled (tokenizer_use_fast = true).
Padding token </s> used as the padding token.
Micro-batch size 2 samples per device.
Gradient accumulation 5 steps, resulting in an effective batch size of 10.
Learning rate $1\times 10^{-5}$.
Training epochs 1 epoch.
Maximum sequence length 5,000 tokens.
Optimizer lion_8bit.
Learning rate schedule constant_with_warmup.
Weight decay 0.01.
Warmup ratio 0.05.
Mixed precision bf16 (automatic).
Gradient checkpointing Enabled to reduce memory usage.
Dataset GSM8K math reasoning dataset.
Dataset configuration name=main, split=train, type=alpaca.
Validation split 5% of the training data.
Logging frequency Every training step.
Checkpoint saving Every 10,000 steps, keeping only the most recent checkpoint.
Reporting Disabled (report_to = none).
Table 11: Mapping from task IDs used in Table 1 to full SuperNaturalInstructions task names and descriptions.
Task ID Full Task Name Description
task108 contextualabusedetection_classification Classify whether a given text contains contextual or implicit abusive language.
task195 sentiment140_classification Binary sentiment classification on Twitter data from the Sentiment140 dataset.
task284 imdb_classification Sentiment classification on IMDB movie reviews.
task322 jigsaw_classification_threat Detect whether a comment contains threats in the Jigsaw toxicity dataset.
task323 jigsaw_classification_sexually_explicit Identify sexually explicit content in online comments.
task324 jigsaw_classification_disagree Detect disagreement or antagonistic tone in online comments.
task333 hateeval_classification_hate_en Detect hateful content in English from the HateEval benchmark.
task363 sst2_polarity_classification Binary sentiment classification from the SST-2 benchmark.
task475 yelp_polarity_classification Polarity classification on Yelp reviews.
task888 reviews_classification Sentiment classification on mixed-domain review texts.
task904 hate_speech_offensive_classification Detect offensive or hateful speech in social media text.
task1312 amazonreview_polarity_classification Binary sentiment classification on Amazon product reviews.
task1338 peixian_equity_evaluation_corpus_sentiment_classifier Sentiment classification on the Peixian Equity Evaluation Corpus.
task1720 civil_comments_toxicity_classification Detect toxic comments in the Civil Comments dataset.
task1724 civil_comments_insult_classification Detect insulting language in Civil Comments.
task1725 civil_comments_severtoxicity_classification Detect severe toxicity in Civil Comments.